A closed cycle gas turbine unit operating is maximum and minimum temperature of 760 and 20 degree Celsius has a pressure ratio of 7/1. Calculate the ideal cycle efficiency and the work ratio
1
Expert's answer
2020-01-15T07:10:09-0500

we can write

"\frac{T_2}{T_1} =\frac{\frac {p_2}{p_1}×(r-1)} {r} (1)"

Where "\frac {p_2}{p_1}=7" , r=1.4

Using (1) we get

"\frac{T_2}{T_1} =1.74"

In our case, we have T1=293 K

"T_2=T_1\u00d71.74=509.8 K"

"\frac{T_2}{T_1} =\frac{T_3}{T_4} (2)"

In our case, we have T3=1033 K

"T_4=\frac{1033}{1.74} =593.7 K"

the ideal cycle efficiency is equal to

"k=1-\frac{T_4-T_1}{T_3-T_2} =43%"


work ratio is equal to

"work ratio=\frac{net work}{gross work} (3)"


net work is equal to

"net work=c_p\u00d7(T_3-T_2)-c_p\u00d7(T_4-T_1)=223.5"


gross work is equal to

"gross work= c_p\u00d7(T_3-T_4)=441"


work ratio=0.5


Answer

ideal cycle efficiency 43%

work ratio=0.5